A man throws balls with the same speed vertically upwards one after the other at an interval of 2seconds.What should be the speed of throw so that more than two balls are in the sky at any time?

Grade:11

5 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

Dear chakrala mounica

the first has to be above the ground when the second ball is thrown from the ground

 

so the first ball has to be above from the ground for 2 seonds.

 

s=ut-1/2*gt^2        t=2  and s=0 when first ball reaches the ground

 

0= 2*u- 1/2*g*4

 

u=g m/s        take g=9.8 m/s^2

u= 9.8 m/s

 

speed has to be more than 9.8 m/s

 

 



Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar..



Askiitians Expert

Rathod Shankar Singh

IIT Bombay

 

Raunak Chaudhury
6 years ago
S=ut-1/2gt×t S=2u-19.6v^2=u^2-2aS0^2=u^2-2×9.8×(2u-19.6)0=u^2- 39.2u - (19.6)^20=(u- 19.6)^2u=19.6 m/secMore than 2 balls speed is greater than 19.6 m/sec
sahil
6 years ago
more than 19.6 m/s Let the required speed of through be u m/s. The time taken to reach maximum height t=u/g ; For two balls to remain in air at any time, t must be greater than 2. ; ∴ u/g >2 ⇒ u > 2(g) ; u > 19.6 m/s
vansh tayal
5 years ago

For at least 3 balls to remain above the ground, the time of flight of the first ball must be more than or equal to 4 s. Because in t = 4 s the third ball leaves the ground.

Now, for the first ball,

Initial velocity = u

Final velocity = -u

Acceleration = -g

So, -u = u – gt

=> u = gt/2

=> u = 9.8 × 4/2

=> u = 19.6 m/s

so the speed should be greater than 19.6 m/s to keep more than 2 balls in the sky at any time.

Thus, the ball must be thrown at speed more than or equal to 19.6 m/s u

Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question below.
 
More than two balls in air means atleast 3 balls need to remain above the ground
Hence, the time of flight of the first ball must be more than or equal to 4 s. Because in t = 4 s the third ball leaves the ground.
Now, s = ut – ½gt2
0 = u(4) – ½ x 9.81 x (4)2
u = 19.62 m/s
Hence, the throwing speed must be greater than 19.62 m/s
 
Hope it helps.
Thanks for the question.
Regards,
Kushagra